# Exact predicates, exact constructions and combinatorics

### v1

@@ Abstract.
This article introduces a general mesh intersection algorithm that exactly computes the so-called Weiler model and that uses it to implement boolean operations with arbitrary multi-operand expressions, CSG (constructive solid geometry) and some mesh repair operations. From an input polygon soup, the algorithm first computes the co-refinement, with an exact representation of the intersection points. Then, the decomposition of 3D space into volumetric regions (Weiler model) is constructed, by sorting the facets around the non-manifold intersection edges (radial sort), using specialized exact predicates. Finally, based on the input boolean expression, the triangular facets that belong to the boundary of the result are classified. This is, to our knowledge, the first algorithm that computes an exact Weiler model. To implement all the involved predicates and constructions, two geometric kernels are proposed, tested and discussed (arithmetic expansions and multi-precision floating-point). As a guiding principle,the combinatorial information shared between each step is kept as simple as possible. It is made possible by treating all the particular cases in the kernel. In particular, triangles with intersections are remeshed using the (uniquely defined) Constrained Delaunay Triangulation, with symbolic perturbations to disambiguate configurations with co-cyclic points. It makes it easy to discard the duplicated triangles that appear when remeshing overlapping facets. The method is tested and compared with previous work, on the existing “thingi10K” dataset (to test co-refinement and mesh repair) and on a new “thingiCSG” dataset made publicly available (to test the full CSG pipeline) on a variety of interesting examples featuring different types of “pathologies”.
@@ 1.1. Why is mesh intersection so hard ?
Once the intersecting triangles are determined, one needs to compute the actual intersection. In general, the coordinates of the intersection between triangles can be expressed as rational fractions (ratios of polynomials in the initial mesh coordinates). Again, in general, they cannot be exactly represented by floating-point numbers. It has several consequences: first, imagine you know the exact coordinates of all the vertices of your mesh, when you convert them into floating point numbers, they will move a little bit. If you do not take care, it may generate new intersections ! The problem of constraining the intersection points to have floating-point coordinates is referred to as “snap rounding”. It requires special care, that is, one needs to ensure the output with floating-point coordinates does not have new intersection and is topologically equivalent (in a certain sense) to the exact result (Devillers et al., 2018). Second, during the computation, one may need to query geometric predicates on the points resulting from intersections, and the answers of these predicates need to be coherent with all the rest ! For instance, consider a triangle $t_{1}$ that has some intersections with other triangles $t_{2},\ldots,t_{n}$. These intersections create segments $e_{1}$ and $e_{2}$ in $t_{1}$, and one needs to re-triangulate $t_{1}$ in a way that conforms with these segments. In other words, one needs to compute a 2D constrained triangulation in $t_{1}$. This 2D constrained triangulation depends on a set of predicates, and these predicates need to be coherent with all the rest. In fact, the situation is even more complicated: as shown in Figure 1-A, consider that $t_{1}$ is a triangle that have intersections with two other triangles $t_{2}$ and $t_{3}$, mutually intersecting. The mesh resulting from the intersection of the three triangles is shown in Figure 1-B. When remeshing $t_{1}$, there will be two segments $e_{1}$ and $e_{2}$ (highlighted in Figure 1-C), (one that corresponds to $t_{1}\cap t_{2}$ and the other one to $t_{1}\cap t_{3}$, that create a new intersection $I$). Now think of what you have to do to compute this intersection: the extremities of $e_{1}$ and $e_{2}$ are intersections between input mesh triangles, and we need to compute their intersection. In this specific case, one can compute instead the intersection between the supporting planes of the three triangles (that solely depend on the input points), as will be explained later, but one needs to keep in mind that the intersection $I$ is a constructed points with coordinates that are rational fractions, that will be later passed through the geometric predicates when computing the constrained triangulation of $t_{1}$. Predicates that depend on constructed points are also used in the subsequent steps of the algorithm, such as the “radial sort” operation involved in the construction of the Weiler model, that needs new predicates, as shall be seen later.
@@
Then it would be tempting to always keep all the combinatorial information. However, doing so has the hidden cost of making the architecture more complicated and more difficult to test. There is a tradeoff to find between a set of independant and easty-to-test pipeline stages connected by a simple communication protocol, or a fully interconnected set of pipeline stages, (slightly) more efficient, but (considerably) harder to design, to test and to debug.
@@ 1.2. Summary of this article’s contributions
This article introduces the first algorithm that computes the so-called Weiler model (Weiler, 1988) exactly, that is, a data structure that stores the decomposition of 3D space into volumetric regions yielded by a set of (possibily intersecting) triangulated surfaces.
The algorithm is based on several components:
  • a multi-thread friendly constrained Delaunay triangulation;
@@
  • an algorithm to simplify the triangulation of co-planar regions.
To implement the required exact predicates and exact constructions, two geometric kernel are described and analyzed:
  • one is based on arithmetric expansions, like in the approach proposed by Shewchuk for predicates (Shewchuk, 1996a), with the difference that arithmetic expansions are also used to store constructed points;
  • the other one is based on multi-precision arithmetics.
For both geometric kernels, I explain how to efficiently implement the arithmetic filters, the predicates, and the symbolic perturbations that ensure the uniqueness of the constrained Delaunay triangulation.
@@
  • The Tingi10K database (Zhou and Jacobson, 2016), used to evaluate the co-refinement algorithm;
  • A new ThingiCSG database with a collection of CSG trees, and the skeleton of an OpenSCAD-compatible CSG engine that can be used to test and benchmark future works111see links at the end of this article.
@@ 1.3. Previous work on mesh intersection
It is also possible to use interval arithmetics, that is, implementing addition, subtraction and product for low,high pair of floating point numbers. Each time 0 is not contained in the [low,high] interval, the sign of the represented number is known. In all other cases, one needs to relaunch the computation with expansions. Since it does not happen often, there is a significant performance gain. Another possibility to quickly determine the sign in the easy configurations is to use arithmetic filters (Meyer and Pion, 2008). The idea shares some similarities with interval arithmetics, with the difference that it computes with the estimated number an error bound, using a combination of static information (deduced from the algebraic expression to be computed) and dynamic information (computed from the actual values passed to the expression). The Predicate Construction Kit (Lévy, 2016) takes an algebraic expression of the predicate, and generates the filter with FPG (Meyer and Pion, 2008), the code that computes it with expansions when the filter fails, as well as symbolic perturbations (Edelsbrunner and Mücke, 1994) for the degenerate configurations (such as 4 cocyclic points in a Delaunay triangulation). It was used to generate the non-standard predicates required to compute the intersection between a Voronoi or a power diagram and a surface or a volume embedded in $n$ dimensions, used in semi-dicrete optimal transport (Lévy, 2015) and its applications in fluid simulation (Lévy, 2022) and cosmology (Lévy et al., 2021; von Hausegger et al., 2022; Nikakhtar et al., 2022).
@@
Using these technologies, algorithms and softwares were developed for robust tetrahedral meshing of triangle soups (that may have intersections), first in the TetWild package (Hu et al., 2018) that was improved and made more efficient a while later, based on a more sophisticated algorithm, dubbed as FastTetWild (Hu et al., 2020). In the same period of time, the Thingi10K dataset was published (Zhou and Jacobson, 2016). It provides the research community with a large database of meshes, with many different cases of degeneracies and inconsistencies. It represents an excellent stress test for mesh intersection algorithms. With objectives and ideas similar to the Predicate Construction Kit (Lévy, 2016) a system for indirect predicates was introduced (Attene, 2020), with in addition the idea of storing intermediate construction (that is, new points generated during the execution of the algorithm). Based on this system, a series of significant advances was published, first to compute mesh arrangements (Cherchi et al., 2020) then boolean operations (Cherchi et al., 2022). Both methods need to compute constrained triangulations. For that, a highly efficient (linear complexity) ear cutting algorithm was proposed (Livesu et al., 2022). The idea of indirect predicates was also used to implement 3D constrained Delaunay triangulation (Diazzi et al., 2023), based on an exact representation of the Steiner points.
To help structuring the information through the processing pipeline, several combinatorial data structures were proposed, such as the Weiler model (Weiler, 1985). This data structure is popular in computational geosciences, where it is used to represent the volumetric relations between rock layers and geologic faults (Mallet, 2002; Sword, 1996; Caumon et al., 2003; Pellerin et al., 2017; Legentil et al., 2022). It constructs relations between volumes from intersecting surfaces by sorting triangles around non-manifold intersection edges. We shall see later in this article how this operation can be exactly implemented. To study the combinatorial aspects of the algorithms, I shall use the notations of combinatorial maps (Lienhardt, 1988). Based on an algebraic specification, the notion of combinatorial maps is well suited to mathematicaly define the operations of a 3D modeler (Bertrand and Dufourd, 1994), as well as multiresolution modeling operators (Kraemer et al., 2007) and some operations involved in hexahedral mesh generation (Lyon et al., 2016). In our context, it facilitates describing and implementing the algorithms in a readable and compact way.
@@
The OpenSCAD software (Wolf and contributors, 2010) is a scriptable CAD package widely used in the fabrication community. It is based on the CGAL computational geometry library (The CGAL Project, 2023a), more precisely, it uses the implementation of Nef complexes (Granados et al., 2003) available in CGAL. To improve computation time, an alternative algorithm, based on co-refinement, was implemented, also in CGAL (Katrioplas et al., 2017). It is significantly faster than the initial implementation (10x to 50x). All the approaches mentioned above follow the exact geometry paradigm (exact predicates and exact constructions, except for the final conversion to floating-point coordinates / snap rounding). There exists a worth mentioning original and highly effective alternative to this paradigm: instead of doing all the computations in exact mode then converting into floating-point coordinates while preserving some topological properties, why not doing all computations in floating point coordinates, while ensuring that each individual computation preserves some properties ? It is the strategy proposed in (Smith and Dodgson, 2006), that builds a network of predicates and operations of increasing dimensions on top of 1D axis-aligned comparisons (that can be done exactly), while ensuring that the mesh stays manifold at each level. A fine-tuned parallel implementation was developed (Lalish, 2022) and integrated as an alternative kernel for OpenSCAD. Since it solely uses single-precision floating point arithmetics, it can also run on the GPU. The performance gain is spectacular (three orders of magnitude). In this article, I follow the more classical approach (exact predicates and exact constructions). I implemented an Open-SCAD compatible software to compare the results with the three geometric kernels mentioned above (§3.2 p.3.2).
@@ 1.4. Overview of this article
The approach presented in this article shares some similarities with the works mentioned above, in particular (Cherchi et al., 2020, 2022), with the following differences:
@@
    exact constructions are used instead of Predicate Construction Kit or indirect predicates;
@@
    the algorithm computes a constrained Delaunay triangulation (CDT), which ensures the uniquenes of the obtained triangulation. The uniqueness of the triangulation is useful to make sure that the same triangles will be generated in configurations with co-planar overlapping facets. Then duplicated triangles can be easily eliminated, whereas previous works need to use an auxilliary data structure to identify overlapping zones. Note that CDT uses the in_circle predicates, that depends on 4 points. Using Predicate Construction Kit or indirect predicates would generate a large number of instances of the in_circle predicates (depending on whether the points are initial vertices or intersections). Here this “combinatorial explosion” is avoided thanks to the exact constructions. In a certain sense, the complexity is pushed towards the kernel;
@@
    the algorithm computes the 3D partition of space into volumes, represented in a data structure called the Weiler model (Weiler, 1985). I shall explain later how this data structure can be efficiently constructed, and how most of the calls to the predicates can be avoided by exploiting the combinatorics. Then I shall use this data structure to implement boolean operations and CSG primitives, still exploiting the combinatorics as much as possible. To my knowledge, this is the first algorithm that exactly computes the Weiler model.
This article introduces an algorithmic pipeline depicted in Figure 2. Given an input polygon soup, the algorithm computes the co-refinement (§2.1 p.2.1), based on an exact constrained Delaunay triangulation constructed in each facet that has intersections (§2.1.3 p.2.1.3). Then, the algorithm creates the Weiler model (§2.2.1 p.2.2.1), that represents all the radial relations around the non-manifold intersection edges as well as the volumetric relations between the regions delimited by the surfaces. Finally, from the input boolean expression, a classification algorithm selects the triangular facets that belong to the boundary of the result (§2.2.3 p.2.2.3). In addition, an optional step merges and re-triangulates co-planar facets (§2.2.4 p.2.2.4). To exactly represent the intersection points, I implemented and tested two different arithmetic kernels (§2.4 p.2.4). One is based on arithmetic expansions (§2.4.1 p.2.4.1), and the other one on multi-precision floating point arithmetics (§2.4.2 p.2.4.2). Both are pre-filtered using interval arithmetics. The uniqueness of the Delaunay triangulation is ensured, even in configurations with cocyclic points, by using symbolic perturbations. This ensures that regions overlapped by multiple coplanar input facets are coherently and uniquely triangulated. Finally, I tested the method and compared with previous work (§3 p.3), on the existing “thingi10K” dataset for the co-refinement algorithm, (§3.1 p.3.1) and on a new “thingiCSG” dataset together with an OpenSCAD compatible CSG engine, both made publicly available333see links at the end of this article, to test the full CSG pipeline on a variety of interesting examples featuring different types of “pathologies” (§3.2 p.3.2).

### v2

@@ Abstract.
This article introduces a general mesh intersection algorithm that exactly computes the so-called Weiler model (also called a 3D arrangement) and that uses it to implement boolean operations with arbitrary multi-operand expressions, CSG (constructive solid geometry) and some mesh repair operations. From an input polygon soup, the algorithm first computes the co-refinement, with an exact representation of the intersection points. Then, the decomposition of 3D space into volumetric regions (Weiler model) is constructed, by sorting the facets around the non-manifold intersection edges (radial sort), using specialized exact predicates. Finally, based on the input boolean expression, the triangular facets that belong to the boundary of the result are classified. The main contribution is a 2D Constrained Delaunay Triangulation with exact coordinates that represent the intersections, thanks to two geometric kernels that are proposed, tested and discussed (arithmetic expansions and multi-precision floating-point). As a guiding principle, the combinatorial information shared between each step is kept as simple as possible. It is made possible by treating all the particular cases in the kernel. In particular, triangles with intersections are remeshed using the (uniquely defined) Constrained Delaunay Triangulation, with symbolic perturbations to disambiguate configurations with co-cyclic points. It makes it easy to discard the duplicated triangles that appear when remeshing overlapping facets. The method is tested and compared with previous work, on the existing “thingi10K” dataset (to test co-refinement and mesh repair) and on a new “thingiCSG” dataset made publicly available111 as well as the main algorithm and the arithmetic kernel based on expansions, see links at the end of the article. (to test the full CSG pipeline) on a variety of interesting examples featuring different types of “pathologies”.
@@ 1.1. Why is mesh intersection so hard?
Once the intersecting triangles are determined, one needs to compute the actual intersection. In general, the coordinates of the intersection between triangles can be expressed as rational fractions (ratios of polynomials in the initial mesh coordinates). Again, in general, they cannot be exactly represented by floating-point numbers. It has several consequences: first, imagine you know the exact coordinates of all the vertices of your mesh, when you convert them into floating point numbers, they will move a little bit. If you do not take care, it may generate new intersections! The problem of constraining the intersection points to have floating-point coordinates is referred to as “snap rounding”. It requires special care, that is, one needs to ensure the output with floating-point coordinates does not have new intersection and is topologically equivalent (in a certain sense) to the exact result (Devillers et al., 2018). Second, during the computation, one may need to query geometric predicates on the points resulting from intersections, and the answers of these predicates need to be coherent with all the rest! For instance, consider a triangle $t_{1}$ that has some intersections with other triangles $t_{2},\ldots,t_{n}$. These intersections create segments $e_{1}$ and $e_{2}$ in $t_{1}$, and one needs to re-triangulate $t_{1}$ in a way that conforms with these segments. In other words, one needs to compute a 2D constrained triangulation in $t_{1}$. This 2D constrained triangulation depends on a set of predicates, and these predicates need to be coherent with all the rest. In fact, the situation is even more complicated: as shown in Figure 1-A, consider that $t_{1}$ is a triangle that has intersections with two other triangles $t_{2}$ and $t_{3}$, mutually intersecting. The mesh resulting from the intersection of the three triangles is shown in Figure 1-B. When remeshing $t_{1}$, there will be two segments $e_{1}$ and $e_{2}$ (highlighted in Figure 1-C), (one that corresponds to $t_{1}\cap t_{2}$ and the other one to $t_{1}\cap t_{3}$, that create a new intersection $I$). Now think of what you have to do to compute this intersection: the extremities of $e_{1}$ and $e_{2}$ are intersections between input mesh triangles, and we need to compute their intersection. In this specific case, one can compute instead the intersection between the supporting planes of the three triangles (that solely depend on the input points), as will be explained later, but one needs to keep in mind that the intersection $I$ is a constructed points with coordinates that are rational fractions, that will be later passed through the geometric predicates when computing the constrained triangulation of $t_{1}$. Predicates that depend on constructed points are also used in the subsequent steps of the algorithm, such as the “radial sort” operation involved in the construction of the Weiler model, that needs new predicates, as shall be seen later.
@@
Then it would be tempting to always keep all the combinatorial information. However, doing so has the hidden cost of making the architecture more complicated and more difficult to test. There is a tradeoff to find between a set of independant and easy-to-test pipeline stages connected by a simple communication protocol, or a fully interconnected set of pipeline stages, (slightly) more efficient, but (considerably) harder to design, to test and to debug.
@@ 1.2. Summary of this article’s contributions
This article presents an algorithm that computes the so-called Weiler model (Weiler, 1988) (also called 3D arrangement) exactly, that is, a data structure that stores the decomposition of 3D space into volumetric regions yielded by a set of (possibily intersecting) triangulated surfaces.
The algorithm is based on several components. The main new contribution presented here is:
  • a new multi-thread friendly constrained Delaunay triangulation algorithm, based on revisiting the flip-based algorithm in a way that minimizes predicates invocations.
To implement the required exact predicates and exact constructions, two new geometric kernel are described and analyzed:
  • one is based on arithmetic expansions, like in the approach proposed by Shewchuk for predicates (Shewchuk, 1996a), with the difference that arithmetic expansions are also used to store constructed points;
  • the other one is based on multi-precision arithmetics.
For both geometric kernels, I explain how to efficiently implement the arithmetic filters, the predicates, and the symbolic perturbations that ensure the uniqueness of the constrained Delaunay triangulation. The main benefit of the set of new algorithms introduced here is a significant performance gain (up to 6x as compared to (Cherchi et al., 2022) and (Zhou et al., 2016)) in degenerate configurations with intersecting co-planar facets (but the integer-based method EMBER (Trettner et al., 2022) remains significantly faster). Degenerate co-planar configurations are present in some of the models in the Thingi10K database, as shown in the Results section. More importantly, such configurations are very often, nearly systematically generated in the CSG trees created py practitioners using modeling tools such as the OpenSCAD language, very common in Thingiverse for instance.
In the article, to help the practitioner who would want to totally re-implement a similar algorithm, I also explain more classical components, also present in the previous work:
@@
  • an algorithm to simplify the triangulation of co-planar regions. This algorithm makes use of the new constrained Delaunay triangulation, efficiently merging co-planar facets, resulting in simpler triangulations. This is especially important when chaining multiple boolean operations present in a deep CSG tree.
@@
  • The Thingi10K database (Zhou and Jacobson, 2016), used to evaluate the co-refinement algorithm;
  • A small set of scanned meshes and CAD meshes with co-planar facets to compare the new method with (Zhou et al., 2016) and (Cherchi et al., 2022);
  • A new ThingiCSG database with a collection of CSG trees, and the skeleton of an OpenSCAD-compatible CSG engine that can be used to test and benchmark future works222see links at the end of the article..
@@ 1.3. Previous work on mesh intersection
It is also possible to use interval arithmetics, that is, implementing addition, subtraction and product for low,high pair of floating point numbers. Each time 0 is not contained in the [low,high] interval, the sign of the represented number is known. In all other cases, one needs to relaunch the computation with expansions. Since it does not happen often, there is a significant performance gain. Another possibility to quickly determine the sign in the easy configurations is to use arithmetic filters (Meyer and Pion, 2008). The idea shares some similarities with interval arithmetics, with the difference that it computes with the estimated number an error bound, using a combination of static information (deduced from the algebraic expression to be computed) and dynamic information (computed from the actual values passed to the expression). The Predicate Construction Kit (Lévy, 2016) takes an algebraic expression of the predicate, and generates the filter with FPG (Meyer and Pion, 2008), the code that computes it with expansions when the filter fails, as well as symbolic perturbations (Edelsbrunner and Mücke, 1994) for the degenerate configurations (such as 4 cocyclic points in a Delaunay triangulation). It was used to generate the non-standard predicates required to compute the intersection between a Voronoi or a power diagram and a surface or a volume embedded in $n$ dimensions, used in semi-dicrete optimal transport (Lévy, 2015) and its applications in fluid simulation (Lévy, 2022) and cosmology (Lévy et al., 2021; von Hausegger et al., 2022; Nikakhtar et al., 2022). With objectives and ideas similar to the Predicate Construction Kit (Lévy, 2016) a system for indirect predicates was introduced (Attene, 2020), with in addition the idea of storing intermediate construction (that is, new points generated during the execution of the algorithm). All the techniques mentioned above are based on floating-point numbers. A possible alternative is to use integer numbers only. In (Trettner et al., 2022; Nehring-Wirxel et al., 2021), plane-based representations are used along with homogenous coordinates. The impact of this carefully designed integer-based representation is a spectacular performance gain without sacrifying robustness.
@@
Early works for computing non-exact mesh intersections were proposed in (Douze et al., 2015), that proposes a fast algorithm for mesh intersections used in the context of multi-view acquisition. The same type of approach was explored in (Barki et al., 2015) with special care for robustness, focusing on the involved geometric predicates. Using this type of approaches, as well as the techniques for robust predicates mentioned in the previous paragraph, algorithms and softwares were developed for robust tetrahedral meshing of triangle soups (that may have intersections), first in the TetWild package (Hu et al., 2018) that was improved and made more efficient a while later, based on a more sophisticated algorithm, dubbed as FastTetWild (Hu et al., 2020). In the same period of time, the Thingi10K dataset was published (Zhou and Jacobson, 2016). It provides the research community with a large database of meshes, with many different cases of degeneracies and inconsistencies. It represents an excellent stress test for mesh intersection algorithms. For instance, it was used the same year to test and validate a robust algorithm to compute the 3D arrangement defined by a set of meshes (Zhou et al., 2016).
Based on the indirect predicates low-level approach mentioned above, a series of significant advances was published, first to compute mesh arrangements (Cherchi et al., 2020) then boolean operations (Cherchi et al., 2022). Both methods need to compute constrained triangulations. For that, a highly efficient (linear complexity) ear cutting algorithm was proposed (Livesu et al., 2022). The idea of indirect predicates was also used to implement 3D constrained Delaunay triangulation (Diazzi et al., 2023), based on an exact representation of the Steiner points.
To help structuring the information through the processing pipeline, several combinatorial data structures were proposed, such as the Weiler model (Weiler, 1985). This data structure is popular in computational geosciences, where it is used to represent the volumetric relations between rock layers and geologic faults (Mallet, 2002; Sword, 1996; Caumon et al., 2003; Pellerin et al., 2017; Legentil et al., 2022). There exists exact algorithms for constructing a Weiler model (or 3D mesh arrangement) (Zhou et al., 2016; Cherchi et al., 2020, 2022) and variants that operate on isosurfaces in a volume (Du et al., 2022). They all follow a very similar pipeline: they first construct relations between volumes from intersecting surfaces by sorting triangles around non-manifold intersection edges, then reconstruct the boundaries volumetric zones. To study the combinatorial aspects of the algorithms, I shall use the notations of combinatorial maps (Lienhardt, 1988). Based on an algebraic specification, the notion of combinatorial maps is well suited to mathematicaly define the operations of a 3D modeler (Bertrand and Dufourd, 1994), as well as multiresolution modeling operators (Kraemer et al., 2007) and some operations involved in hexahedral mesh generation (Lyon et al., 2016). In our context, it facilitates describing and implementing the algorithms in a readable and compact way.
@@
The OpenSCAD software (Wolf and contributors, 2010) is a scriptable CAD package widely used in the fabrication community. It is based on the CGAL computational geometry library (The CGAL Project, 2023a), more precisely, it uses the implementation of Nef complexes (Granados et al., 2003) available in CGAL. To improve computation time, an alternative algorithm, based on co-refinement, was implemented, also in CGAL (Katrioplas et al., 2017). It is significantly faster than the initial implementation (10x to 50x). All the approaches mentioned above follow the exact geometry paradigm (exact predicates and exact constructions, except for the final conversion to floating-point coordinates / snap rounding). There exists a worth mentioning original and highly effective alternative to this paradigm: instead of doing all the computations in exact mode then converting into floating-point coordinates while preserving some topological properties, why not doing all computations in floating point coordinates, while ensuring that each individual computation preserves some properties? It is the strategy proposed in (Smith and Dodgson, 2006), that builds a network of predicates and operations of increasing dimensions on top of 1D axis-aligned comparisons (that can be done exactly), while ensuring that the mesh stays manifold at each level. A fine-tuned parallel implementation was developed (Lalish, 2022) and integrated as an alternative kernel for OpenSCAD. Since it solely uses single-precision floating point arithmetics, it can also run on the GPU. The performance gain is spectacular (three orders of magnitude). In this article, I follow the more classical approach (exact predicates and exact constructions). I implemented an Open-SCAD compatible software to compare the results with the three geometric kernels mentioned above (§3.3 p.3.3).
@@ 1.4. Overview of this article
The approach presented in this article shares some similarities with the works mentioned above, in particular (Cherchi et al., 2020, 2022; Zhou et al., 2016), with the following differences:
@@
    unlike in (Cherchi et al., 2020, 2022) exact constructions are used instead of Predicate Construction Kit or indirect predicates. Note that exact constructions (from CGAL) are also used in (Zhou et al., 2016). We further investigate the geometric kernel with two new numbers representations and the associated predicates, for significant gains (up to 6x) in degenerate configurations, as compared to the previous works mentioned above;
@@
    the algorithm computes a constrained Delaunay triangulation (CDT), which ensures the uniquenes of the obtained triangulation. This comes as a higher computational cost, related to the in_circle() predicate, as compared to (Cherchi et al., 2020) that computes a (non-Delaunay) constrained triangulation. The uniqueness of the triangulation is useful to make sure that the same triangles will be generated in configurations with co-planar overlapping facets. Then duplicated triangles can be easily eliminated, whereas previous works need to use an auxilliary data structure to identify overlapping zones. Note that the in_circle predicates depends on 4 points, that can have different expressions (they can be input data points, or intersections of different types). Using Predicate Construction Kit or indirect predicates would generate a large number of instances of the in_circle predicates (depending on whether the points are initial vertices or intersections). Here this “combinatorial explosion” is avoided thanks to the exact constructions. In a certain sense, the complexity is pushed towards the kernel. Another benefit of this efficient CDT algorithm is the ability of efficiently re-triangulating zones made of co-planar triangles. This is expecially important when chaining boolean operations, as in the evaluation of CSG trees, to avoid generating a huge number of triangles;
@@
    the algorithm computes a 3D partition of space into volumes, that is, a 3D mesh arrangement, represented in a data structure called the Weiler model (Weiler, 1985). I shall explain later how this data structure can be efficiently constructed, and how most of the calls to the predicates can be avoided by exploiting the combinatorics. Then I shall use this data structure to implement boolean operations and CSG primitives, still exploiting the combinatorics as much as possible. This is very similar to the 3D mesh arrangement construction in (Zhou et al., 2016). The difference is that I describe it using the notion of combinatorial map (Lienhardt, 1994), that yields a compact form of the combinatorial operations such as the classification.
This article introduces an algorithmic pipeline depicted in Figure 2. Given an input polygon soup, the algorithm computes the co-refinement (§2.1 p.2.1), based on an exact constrained Delaunay triangulation constructed in each facet that has intersections (§2.1.3 p.2.1.3). Then, the algorithm creates the Weiler model (§2.2.1 p.2.2.1), that represents all the radial relations around the non-manifold intersection edges as well as the volumetric relations between the regions delimited by the surfaces. Finally, from the input boolean expression, a classification algorithm selects the triangular facets that belong to the boundary of the result (§2.2.3 p.2.2.3). In addition, an optional step merges and re-triangulates co-planar facets (§2.2.4 p.2.2.4). To exactly represent the intersection points, I implemented and tested two different arithmetic kernels (§2.4 p.2.4). One is based on arithmetic expansions (§2.4.1 p.2.4.1), and the other one on multi-precision floating point arithmetics (§2.4.2 p.2.4.2). Both are pre-filtered using interval arithmetics. The uniqueness of the Delaunay triangulation is ensured, even in configurations with cocyclic points, by using symbolic perturbations. This ensures that regions overlapped by multiple coplanar input facets are coherently and uniquely triangulated. Finally, I tested the method and compared with previous work (§3 p.3), on the existing “thingi10K” dataset for the co-refinement algorithm, (§3.1 p.3.1) and on a new “thingiCSG” dataset together with an OpenSCAD compatible CSG engine, both made publicly available444both publicly available, as well as the main algorithm and the arithmetic kernel based on expansions, see links at the end of the article., to test the full CSG pipeline on a variety of interesting examples featuring different types of “pathologies” (§3.3 p.3.3).